OpenStudy (anonymous):

Can you apply L'Hopital's Rule multiple times? Also, help with finding the limit of (3x)^(1/3x) algebraically please?

OpenStudy (kc_kennylau):

as long as it's still \(\dfrac00\) or \(\dfrac\infty\infty\) you can use l'hopital rule as many times as you want

OpenStudy (kc_kennylau):

and what does x tend to?

OpenStudy (anonymous):

Towards infinity again

OpenStudy (kc_kennylau):

you mean \(\Large\lim\limits_{x\rightarrow\infty}(3x)^{\frac13x}\)?

OpenStudy (kc_kennylau):

or \(\Large\lim\limits_{x\rightarrow\infty}(3x)^{\frac1{3x}}\)?

OpenStudy (anonymous):

The second one, sorry for not clearing that up earlier

OpenStudy (kc_kennylau):

not your fault :)

OpenStudy (kc_kennylau):

try taking the log of that function :)

OpenStudy (kc_kennylau):

remember \(\large x=e^{\ln x}\)

OpenStudy (anonymous):

I'm not quite sure how to do that

OpenStudy (kc_kennylau):

ok i'll do the first step for you :)

OpenStudy (kc_kennylau):

\[\Large\begin{array}{l}&\lim_{x\rightarrow\infty}(3x)^{\frac1{3x}}\\=&e^{\lim\limits_{x\rightarrow\infty}\ln(3x)^{\frac1{3x}}}\\=&e^{\lim\limits_{x\rightarrow\infty}\frac1{3x}\ln(3x)}\end{array}\]

OpenStudy (kc_kennylau):

now you can apply the l'hopital rule :)

OpenStudy (anonymous):

How would I apply it in this case?